OpenStudy (anonymous):

a triangle has side lengths 15,23 and 28 is it a right triangle??? explain A triangle has side lengths 16,63 and 65 is it a right triangle? explain

OpenStudy (raden):

just prove, does the value of 15^2 + 23^2 = 28^2 ? LHS = or not = with RHS ? do same idea, for 16, 63, 65 triangle

OpenStudy (anonymous):

thankd